Do 70uA test for MCC to see if they can transport the beam; they'll keep the tune and continue to send beam in order to tune for the Compton. When the target is cool we'll try to establish 70uA on PbD#1. We had ramped the magnets down until the target cool down is complete, so in addition to all the normal stuff they will have to be turned back on - but Jack wants to be the one to ramp the right dipole.
If we can't run with 70uA we'll do 50uA, as long as we have both spectrometer arms. If the right dipole goes or for whatever reason we don't have both arms and we can't do 70uA we'll go back to PbD#3 to check the asymptotic temperatures (run for a couple of hours, see log entry 326933).
normal items (once target is cooled):
1) call Jack to ramp the right dipole - ask about other magnets (once target is cooled) 2) turn the other L and R HRS magnets on if Jack okays it 3) make sure charge feedback and dithering are on 4) make sure fast feedback is on 5) watch raster synch
special items (if at 70 uA):
1) watch the collimator temperatures (need to be less than 600 C, or 6 V) 2) watch target vacuum 3) do periodic spot++ using T5 and 5 uA to see PbD#1 density
*** Call Juliette or Bob to establish 70uA on the PbD#1 target - you can do the normal beam recovery procedure and go up to 50uA as usual until they arrive.***